OpenStudy (anonymous):

How do I get the lowest common multiple with: 2233 and 333?

OpenStudy (anonymous):

How do I get the lowest common multiple of 36, 27. Ignore first question...

OpenStudy (anonymous):

you find a number that will divide evenly into both of your numbers

OpenStudy (anonymous):

I know that...

OpenStudy (anonymous):

The first question was more fun - if a little weird, considering they were coprime.

OpenStudy (anonymous):

The answer is 3

OpenStudy (anonymous):

The answer isn't 3 ^

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

my bad i was thinking about something else geesh

OpenStudy (anonymous):

36 = 4 * 9 27 = 3 * 9 Does that help?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

36 = 4 X 9 27 = 3 X 9

OpenStudy (anonymous):

To find the lowest common multiple, you can find their highest common factor first. i.e. if n = 5 * k and m = 7 * k , their lowest common multiple is 5 * 7 * k Or you can write out their multiples until one is the same.

OpenStudy (anonymous):

????????

OpenStudy (anonymous):

????

OpenStudy (anonymous):

newton

OpenStudy (anonymous):

The LCM is 108

OpenStudy (anonymous):

The lowest common multiple is 3 x 4 x 9

OpenStudy (anonymous):

plz help me write my essay

OpenStudy (anonymous):

inewton

OpenStudy (anonymous):

I don't know k and m.

OpenStudy (anonymous):

its 108

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i newtown

OpenStudy (anonymous):

It was just an example. I was trying to help you learn a method yo can apply again and not tell you the answer.

OpenStudy (anonymous):

newtom

OpenStudy (anonymous):

Mary I'm not going to write your essay, now go away.

OpenStudy (anonymous):

ok but plz

OpenStudy (anonymous):

The lcm is 108

OpenStudy (anonymous):

Why ON EARTH would I write your essay.

OpenStudy (anonymous):

108 = 3 x 4 x 9 ....

OpenStudy (anonymous):

that helps me its due tommorow

OpenStudy (anonymous):

He said no so NO.

OpenStudy (anonymous):

model chic how did you figure it out?

OpenStudy (anonymous):

I wrote out all the multiples

OpenStudy (anonymous):

Thank you all...

OpenStudy (anonymous):

I don't have any more questions...

OpenStudy (anonymous):

Thank you INewton and model_chic13 for helping me...